OpenStudy (anonymous):

Is this anywhere correct? F(t_2-t_1)=m(V_2-V_1) N(s_2-s_1)=(kg)(m/s_2-m/s_1) (kg)((m/s)^2)(s_1)=(kg)(m/s_1) not homogeneous?

OpenStudy (anonymous):

@jim_thompson5910 can you help me? :)

jimthompson5910 (jim_thompson5910):

What are you trying to show again?

OpenStudy (anonymous):

I need to determine whether or not the original equation is dimensionally homogeneous.

jimthompson5910 (jim_thompson5910):

So the original equation is F(t_2-t_1)=m(V_2-V_1)

jimthompson5910 (jim_thompson5910):

correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

t_2 and t_1 are times in seconds so t_2 - t_1 is a time in seconds

jimthompson5910 (jim_thompson5910):

So F(t_2-t_1)=m(V_2-V_1) becomes F(seconds - seconds)=m(V_2-V_1) which becomes F(seconds)=m(V_2-V_1)

jimthompson5910 (jim_thompson5910):

Same idea with V_2 and V_1, but these are velocities in m/s V_2-V_1 m/s - m/s m/s Note: m and s are NOT variables, they are units, so m/s is NOT a variable...it's a unit So we get F(seconds) = m(m/s) Do you see how I'm getting all this?

OpenStudy (anonymous):

yes, I understand that. Then, F= (kg)(m/s)^2... So, (kg)(m/s)^2(s)=(kg)(m/s) right?

jimthompson5910 (jim_thompson5910):

Units for force F = Newtons F = kg*(m/s^2) t_2 and t_1 = seconds (s) V_2 and V_1 = m/s m = kg Sadly there are two 'm' terms here, one for meters and one for mass...hopefully that's not too confusing -------------------- F(t_2-t_1)=m(V_2-V_1) F(seconds) = kg(m/s) (kg*(m/s^2))(seconds) = kg(m/s) (kg*(m/s^2))(s) = kg(m/s) kg*m/s = kg*m/s So the dimensions match

jimthompson5910 (jim_thompson5910):

Oops, I meant to write "Units" instead of "units for force", but you get the idea

OpenStudy (anonymous):

(kg*(m/s^2))(s) = kg(m/s) kg*m/s = kg*m/s how'd you go front that to that?

jimthompson5910 (jim_thompson5910):

The s^2 in the denominator and the 's' in the numerator on the left side multiply and one pair of 's' terms cancel So (1/s^2)*(s/1) = 1/s

jimthompson5910 (jim_thompson5910):

That's how (kg*(m/s^2))(s) simplifies to kg*m/s

OpenStudy (anonymous):

oh okay gotcha. and one more I did, but I'm unsure if it's correct... F=mV (kg*(m/s^2))=(kg*(m/s)) not homogeneous?

jimthompson5910 (jim_thompson5910):

no it's not, the left side has a s^2 in the denominator but the right side only has 's' in the denominator so it's not homogeneous

OpenStudy (anonymous):

okay, so yay got one right lol! thanks for the help. greatly appreciated.

jimthompson5910 (jim_thompson5910):

you're welcome